# Device-independent certification of desirable properties with a confidence interval

### v1

@@ V Discussion
Tomography and witnesses are two commonly employed toolkits for certifying the desirable properties of quantum devices [96]. In recent years, as it involves only a minimal set of assumptions, the device-independent (DI) paradigm has offered an appealing alternative to these conventional means. Nonetheless, many DI certification schemes, e.g., [15, 16, 17, 18, 20, 21, 22, 23, 24, 25, 26, 28, 29], rely on access to the ideal quantum correlation $\vec{P}_{\mathcal{Q}}$ (or the actual Bell-inequality violation due to $\vec{P}_{\mathcal{Q}}$). In practice, this is unrealistic for two reasons: (1) we always have access to only a finite amount of experimental data, and (2) actual experimental trials are typically not independent and identically distributed (i.i.d.).
To this end, very specialized tools have been developed for the task of randomness generation, quantum key distributions, and even the self-testing [86, 97, 98] of quantum states. Among them, the possibility of using hypothesis testing (based on the PBR protocol [59]) for self-testing with finite data was first discussed in [86] (see also [97] for a different approach). Meanwhile, it is long known [41, 60, 89] that hypothesis testing in a Bell test can also be carried out using a martingale-based protocol. Here, we demonstrate the viability and versatility of such hypothesis-testing-based approaches for the general problems of DI certification.
Central to our finding is the observation that many desirable quantum properties that one wishes to certify can be characterized by (the complement of) some convex set $\mathcal{C}_{\mathcal{P}}$ in the space of correlation vectors $\{\vec{P}\}$. In other words, if a given $\vec{P}_{\mathcal{Q}}$ lies outside $\mathcal{C}_{\mathcal{P}}$, a Bell-like inequality can be provided to witness this fact. This separating hyperplane then provides the basis for our martingale-based protocol for DI certification. On the other hand, if $\mathcal{C}_{\mathcal{P}}$ itself admits a semidefinite programming characterization like the kind proposed in [65, 66, 67, 16, 21], then the problem of minimizing the statistical distance to $\mathcal{C}_{\mathcal{P}}$ can be cast as a conic program, which can be readily solved using existing solvers, such as $\mathsf{MOSEK}$ [91]. In turn, the PBR protocol provides an optimized Bell-like inequality that facilitates hypothesis testing.
@@
In the i.i.d. setting, the PBR protocol is known to be asymptotically optimal (in terms of its confidence-gain rate). However, we see from Figs. 2, 4 and 6 that for a relatively small number of trials and with the right choice of the Bell function, the martingale-based protocol performs equally well, if not better. A similar observation was also noted in [99] where the authors therein compare the PBR method with the Chernoff-Hoeffding bound in determining the success probability of Bernoulli trials. In our case, this is not surprising as the PBR method does not presuppose a Bell-like inequality but rather sacrifices some of the data to determine one. Indeed, if we equip the PBR protocol with the optimized Bell-like inequality right from the beginning, its performance is, as expected, no worse than the martingale-based protocol. See Figs. 8, 9, 12, 11, 13 and 14 in Appendix A for some explicit examples.
@@
Meanwhile, we also see from Figs. 3, 5 and 7 that for several cases that we have investigated, one’s intuitive choice of the Bell function for the martingale-based method can lead to a relatively poor confidence-gain rate, and hence impairs its efficiency to produce a good $p$-value bound, see Figs. 12 and 14. For example, even though the titled CHSH Bell inequality of Eq. 19 is known to self-test all entangled two-qubit pure states |$\ket{\psi(\theta)}$, this choice of the Bell function in the martingale-based method leads to a worse performance compared with using the CHSH Bell function, which, in turn, gives a suboptimal performance compared with that derived from the PBR protocol, see Fig. 7. At this point, it is worth reiterating that both protocols do not require the assumption that the experimental trials are i.i.d, even though we have only given, for simplicity, examples with i.i.d. trials.
@@
Several research directions naturally follow from the present work. Firstly, even though our hypothesis-testing-based approaches enable rigorous DI certification with a confidence interval, by virtue of the techniques involved, one can only make a rather weak certification: out of the many experimental trials, we can be sure that at least one involves a setup that exhibits the desired property (say, with $99\%$ confidence). This is evidently far from satisfactory. A preferable certification scheme should allow one to comment on the general or average behavior of all the measured samples, as has been achieved in [98, 58] for self-testing. Given that self-testing with a high fidelity is technically challenging, it is still of interest to devise a general recipe for certifying the average behavior of other more specific properties (such as entanglement), which may already be sufficient for the information processing task at hand.
@@
At this point, it is worth noting that if the i.i.d. assumption is somehow granted, then our protocols also certify the quality of the setup for every single runs, including those that have not been measured. In this case, once a sufficiently small $p$-value bound is obtained, one can stop measuring the rest of the systems and use them, instead, for the information processing tasks of interest. Of course, since the i.i.d. assumption is generally not warranted, a protocol that achieves certification for some fraction of the copies while leaving the rest useful for subsequent tasks will be desirable. This has been considered for one-shot distillable entanglement in [19] and the self-testing fidelity in [58]. Again, a general treatment will be more than welcome (see, e.g. [100]).
@@ Acknowledgements.
We are grateful to Yanbao Zhang for many enlightening discussions. This work is supported by the National Science and Technology Council (formerly Ministry of Science and Technology), Taiwan (Grants No. 107-2112-M-006-005-MY2 and 109-2112-M-006-010-MY3, 111-2112-M-005-007-MY4, 112-2119-M-001-004, 112-2119-M-001-006, 112-2124-M-002-003, 112-2628-M-006-007-MY4).
@@ Appendix A Miscellaneous results
In this Appendix, we provide some supplemental results to further illustrate the relative strength of the two hypothesis-testing protocols. For that matter, we extend some of the Figures shown in the main text to include two other plots in each of them. The first of this, dubbed “PBR-simplified” consists of the simplified implementation of the PBR protocol, where we use directly the relative frequency $\vec{f}$, instead of the regularized frequency $\vec{f}_{\text{reg}}$ in Eq. 29 and Eq. 30.
Next, to see the best possible performance that the PBR protocol can offer, we substitute $\vec{f}_{\text{reg}}$ by the ideal quantum correlation $\vec{P}_{\mathcal{Q}}$ used in the generation of the data in Eq. 29 and Eq. 30. In other words, this gets rid of the statistical fluctuations involved in the estimation of $\vec{P}_{\mathcal{Q}}$ right from the beginning. We call this “PBR-ideal” and this offers the best performance that one can ever get from the PBR protocol, as it amounts to using the right optimized Bell-like inequality right from the start.
@@ A.1 Negativity
Clearly, as we can see in Fig. 8, the performance of the PBR-simplified protocol is far worse than the PBR protocol. For example, after $N=10^{5}$ trials, the certifiable negativity is only about the same as that achieved with the PBR protocol with $N=2\times 10^{4}$ trials. Similarly, we see in Fig. 9 that with this protocol, the rate at which the $p$-value bound decreases for hypothesis 1 for $\mathcal{N}_{0}=0.3$ and $0.4$ is far slower than the other protocols.
Figure 9: Plot of -$\log 2p^{\text{(prot)}}$ vs $N$ for hypothesis 1 with $\mathcal{N}_{0}=0.3$ and $0.4$. The parameters are the same as those described in Fig. 2.
On the other hand, these results also illustrate that for the hypothesis testing involved in Fig. 8, the performance of PBR is not so much different from PBR-ideal after all.
In the main text, we mention, in relation to Fig. 3, a possible cause of the noticeable discrepancy in the confidence-gain rates obtained between the two protocols. In Fig. 10, we see that, indeed, for any partially entangled two-qubit state |$\ket{\psi(\theta)}$, the negativity lower bound obtained directly from the CHSH Bell violation, cf.  Eq. 10, is far from tight compared to that obtained by solving Section II.2.1 for, say, level three of the hierarchy introduced in [16]. Interestingly, we also see from Fig. 10 that negativity lower bound continues to improve even at level eleven of the hierarchy. See [101] for closely related studies on the convergence of this and other SDP hierarchies.
Figure 10: Comparison of the actual negativity (top) of $\psi(\theta)$, Eq. 17a, and their various device-independent lower bounds. The computation for the negativity lower bound based on the underlying correlation $\vec{P}_{\theta}$, derived from Section II.2.3, is obtained by solving Section II.2.1 whereas the bottom (dashed) line follows directly from the CHSH Bell inequality violation of these states, as given in Eqs. 10 and 18.

### v2

@@ V Discussion
Tomography and witnesses are two commonly employed toolkits for certifying the desirable properties of quantum devices [99]. In recent years, the device-independent (DI) paradigm has offered an appealing alternative to these conventional means as it involves only a minimal set of assumptions. Nonetheless, many DI certification schemes, e.g., [16, 17, 18, 19, 21, 22, 23, 24, 25, 26, 27, 29, 30], implicitly assumes that the underlying quantum correlation $\vec{P}_{\mathcal{Q}}$ (or the actual Bell-inequality violation due to $\vec{P}_{\mathcal{Q}}$) is known. In practice, this is unrealistic for two reasons: (1) we always have access to only a finite amount of experimental data, and (2) actual experimental trials are typically not independent and identically distributed (i.i.d.).
To this end, very specialized tools have been developed for the task of randomness generation, quantum key distributions, and the self-testing [87, 100, 101] of quantum states. Among them, the possibility of using hypothesis testing (based on the PBR protocol [61]) for self-testing with finite data was first discussed in [87] (see also [100] for a different approach). Meanwhile, it is long known [43, 62, 91] that hypothesis testing in a Bell test can also be carried out using a martingale-based protocol. Here, we demonstrate the viability and versatility of such hypothesis-testing-based approaches for the general problem of DI certification.
Central to our finding is the observation that many desirable quantum properties $\mathcal{P}$ that one wishes to certify can be characterized by (the complement of) some convex set $\mathcal{C}_{\mathcal{P}}$ in the space of correlation vectors $\{\vec{P}\}$. In other words, if a given $\vec{P}_{\mathcal{Q}}$ lies outside $\mathcal{C}_{\mathcal{P}}$, a Bell-like inequality can be provided to witness this fact. This separating hyperplane then provides the basis for our martingale-based protocol for DI certification. On the other hand, if $\mathcal{C}_{\mathcal{P}}$ itself admits a semidefinite programming characterization like the kind proposed in [67, 68, 69, 17, 22], then the problem of minimizing the statistical distance to $\mathcal{C}_{\mathcal{P}}$ can be cast as a conic program, which can be readily solved using existing solvers, such as $\mathsf{MOSEK}$ [93]. In turn, the PBR protocol provides an optimized Bell-like inequality that facilitates the corresponding hypothesis testing.
@@
In the i.i.d. setting, the PBR protocol is known to be asymptotically optimal (in terms of its confidence-gain rate). However, we see from Figs. 2, 4 and 6 that for a relatively small number of trials and with the right choice of the Bell function, the martingale-based protocol performs equally well, if not better. A similar observation was also noted in [102] where the authors therein compare the PBR method with the Chernoff-Hoeffding bound in determining the success probability of Bernoulli trials. In our case, this is not surprising as the PBR method does not presuppose a Bell-like inequality but rather sacrifices some of the data to determine one. Indeed, if we equip the PBR protocol with the optimized Bell-like inequality right from the beginning, its performance is, as expected, no worse than the martingale-based protocol. See Figs. 10, 11, 14, 13, 15 and 16 in Appendix A for some explicit examples.
@@
Meanwhile, we also see from Figs. 3, 5 and 7 that for several cases that we have investigated, one’s intuitive choice of the Bell function for the martingale-based method can lead to a relatively poor confidence-gain rate, and hence impairs its efficiency to produce a good $p$-value bound, see Figs. 14 and 16. For example, even though the titled CHSH Bell inequality of Eq. 19 is known to self-test all entangled two-qubit pure states |$\ket{\psi(\theta)}$, this choice of the Bell function in the martingale-based method leads to a worse performance (for bounding the target-state-fidelity) compared with using the CHSH Bell function, which, in turn, gives a suboptimal performance compared with that derived from the PBR protocol, see Fig. 7. At this point, it is worth reiterating that both protocols do not require the assumption that the experimental trials are i.i.d, even though we have only given, for simplicity, examples with i.i.d. trials.
@@
Several research directions naturally follow from the present work. Firstly, even though our hypothesis-testing-based approaches enable rigorous DI certification with a confidence interval, by virtue of the techniques involved, one can only make a rather weak certification: out of the many experimental trials, we can be sure that at least one involves a setup that exhibits the desired property (say, with $99\%$ confidence). This is evidently far from satisfactory. A preferable certification scheme should allow one to comment on the general or average behavior of all the measured samples, as has been achieved in [101, 60] for self-testing.
Given that self-testing with a high fidelity is technically challenging, it is still of interest to devise a general recipe for certifying the average behavior of other more specific properties (such as entanglement, steerability, etc.), which may already be sufficient for the specific information processing task at hand. However, note that the rejection of a null hypothesis on the average behavior (e.g., average negativity $\overline{\mathcal{N}(\rho)}\leq\mathcal{N}_{0}$) necessarily entails the rejection of the corresponding null hypothesis for all trials (e.g., $\mathcal{N}(\rho)\leq\mathcal{N}_{0}$ in every trial). Thus, we may expect a tradeoff when switching from the current kind of hypothesis testing to that for an average behavior.
@@
Also worth noting that if the i.i.d. assumption is somehow granted, then our protocols also certify the quality of the setup for every single runs, including those that have not been measured. In this case, once a sufficiently small $p$-value bound is obtained, one can stop measuring the rest of the systems and use them, instead, for the information processing tasks of interest. Of course, since the i.i.d. assumption is generally not warranted, a protocol that achieves certification for some fraction of the copies while leaving the rest useful for subsequent tasks will be desirable. This has been considered for one-shot distillable entanglement in [20] and the self-testing fidelity in [60]. Again, a general treatment will be more than welcome (see, e.g. [103]).
@@ Acknowledgements.
We are grateful to Jean-Daniel Bancal, Gelo Tabia, and Yanbao Zhang for many enlightening discussions. This work is supported by the National Science and Technology Council, Taiwan (Grants No. 107-2112-M-006-005-MY2 and 109-2112-M-006-010-MY3, 111-2112-M-005-007-MY4, 112-2119-M-001-004, 112-2119-M-001-006, 112-2124-M-002-003, 112-2628-M-006-007-MY4) and the Foxconn Research Institute, Taipei, Taiwan.
@@ Appendix A Miscellaneous results
In this Appendix, we provide some supplemental results to further illustrate the relative strength of the two hypothesis-testing protocols. For that matter, we extend some of the Figures shown in the main text to include two other plots in each of them. The first of this, dubbed “PBR-simplified” consists of the simplified implementation of the PBR protocol, where we use in Eq. 29 and Eq. 30 the relative frequency $\vec{f}$ computed from the raw data, instead of the regularized frequency $\vec{f}_{\text{reg}}$
On the other hand, to see the best possible performance that one could hope for via any kind of implementation of the PBR protocol, we also replace $\vec{f}_{\text{reg}}$ in Eq. 29 and Eq. 30 by the ideal quantum correlation $\vec{P}_{\mathcal{Q}}$ used in the generation of the data. This gets rid of the statistical fluctuations involved in the estimation of $\vec{P}_{\mathcal{Q}}$ right from the beginning. We call this “PBR-ideal” as it amounts to using the optimal Bell-like inequality right from the beginning.
@@ A.1 Negativity
A.1.1 Certification based on the data generated from $\vec{P}_{\text{CHSH}}$
@@
Clearly, as we can see in Fig. 10, the performance of the PBR-simplified protocol is far worse than the PBR protocol. For example, after $N=10^{5}$ trials, the certifiable negativity is only about the same as that achieved with the PBR protocol with $N=2\times 10^{4}$ trials. Similarly, we see in Fig. 11 that with this protocol, the rate at which the $p$-value bound for Null Hypothesis 1 with $\mathcal{N}_{0}=0.3$ and $0.4$ decreases at a rate that is far slower than the other protocols. On the other hand, we also see from these figures that the performance of PBR-ideal is essentially the same as the one given by the martingale-based protocol, an observation consistent with what one would expect from Fig. 3 (for $\theta=\frac{\pi}{4}$).
Figure 11: Plot of -$\log 2p^{\text{(prot)}}$ vs $N$ for Null Hypothesis 1 with $\mathcal{N}_{0}=0.3$ and $0.4$. The parameters are the same as those described in Fig. 2.
A.1.2 Certification based on the data generated from $\vec{P}_{\theta}$
In the main text, we mention, in relation to Fig. 3, a possible cause of the noticeable discrepancy in the confidence-gain rates obtained between the two protocols. In Fig. 12, we see that, indeed, for any partially entangled two-qubit state |$\ket{\psi(\theta)}$, the negativity lower bound obtained directly from the CHSH Bell violation, cf.  Eq. 10, is far from tight compared to that obtained by solving Eq. 9 for, say, level-3 of the hierarchy introduced in [17]. Interestingly, we also see from Fig. 12 that negativity lower bound continues to improve even at level-12 of the hierarchy. See [104] for closely related studies on the convergence of this and other SDP hierarchies.
Figure 12: Comparison of the actual negativity (top) of $\psi(\theta)$, Eq. 17a, and their various device-independent lower bounds. The computation for the negativity lower bound based on the underlying correlation $\vec{P}_{\theta}$, derived from Eq. 17, is obtained by solving Eq. 9 whereas the bottom (dashed) line follows directly from the CHSH Bell inequality violation of these states, as given in Eqs. 10 and 18.
A.1.3 Certification based on the data generated from $\vec{P}_{\text{CGLMP}}$
@@
Figure 13: Extended plot from Fig. 4 where we also include the results from the simplified implementation of the PBR protocol “PBR-simplified” and the ideal protocol “PBR-ideal”.